You walk 34 m at an angle of 30 degrees N of E. You then walk 45 m at an angle of 30 degrees N of W. Assume E points in the +x direction. Assuming you started from <0,0> where are you? 2.
Answer #1

North 45 m 34 m зо 30 West X1 0 X2 East South

Using a trigonometric identity in RHS, we get

cos 300 = x1 / (34 m)

x1 = [(34 m) (0.8660)]

x1 = 29.4 m

Using a trigonometric identity in LHS, we get

cos 300 = x2 / (45 m)

x2 = [(45 m) (0.8660)]

x2 = 38.9 m

Assuming we started from < 0, 0 >, then our position located at an origin which will be given as :

we know that, d = x1 + x2

d = [(29.4 m) + (38.9 m)]

d = 68.3 m